OpenStudy (anonymous):

4x + 2(x - 3) = 8x + 12 How many solutions ?

OpenStudy (solomonzelman):

\(\large\color{black}{ 4x + 2(x - 3) = 8x + 12 }\) \(\large\color{black}{ 4x + 2x - 6 = 8x + 12 }\) \(\large\color{black}{ 6x - 6 = 8x + 12 }\) \(\large\color{black}{ 6x - 6 = 8x -8+ 20 }\) \(\large\color{black}{ 6(x-1) = 8(x-1)+ 20 }\) \(\large\color{black}{ -2(x-1) = 20 }\) .... you tell me. Sorry for the creativity, as opposed to the shortest approach.